The probability that a person, living in a certain city, owns a dog is estimated to be 0.3. Find the probability that the




tenth person randomly interviewed in that city is the fifth one to own a dog

1
Expert's answer
2022-05-04T14:37:10-0400

We assume that persons are independent. We consider "10" independent random variables "X_i", "i=1,\ldots,10" and each of these variables has Bernoulli distribution. These variables correspond to randomly selected persons. Each variable takes value ('a person does not have a dog') with a probability "0.7" and "1" ('a person has a dog') with a probability "0.3". Suppose that "a_1,a_2,\ldots,a_9" are numbers such that "4" of them are equal to "1" and "5" of them are . The probability is equal to: "p=\sum_{a_1,a_2,\ldots,a_9}P(X_1=a_1,X_2=a_2,\ldots,X_9=a_9,X_{10}=1)", where the sum is taken over all numbers "a_1,a_2,\ldots,a_9" such that "4" of them are equal to "1" and "5" of them are . We point out that "P(X_1=a_1,X_2=a_2,\ldots,X_9=a_9,X_{10}=1)=(0.3)^5(0.7)^5". The sum contains "C_{9}^4=\frac{9!}{5!4!}" terms. Thus, "p=C_9^4(0.3)^5(0.7)^5\approx0.051" (it is rounded to 3 decimal places)

Answer: "0.051"(it is rounded to 3 decimal places)